OpenStudy (anonymous):

Which of the following is the slope between the points (2, 3) and (-3, 0)? - -

OpenStudy (anonymous):

I can't see it???

OpenStudy (anonymous):

- 5/ 3 5/3 -3/5 3/5 options^

OpenStudy (anonymous):

\[m = \frac{ (y2-y1) }{ (x2-x1) }\]

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

hartnn (hartnn):

what you get ?

OpenStudy (anonymous):

no clue, i forgot how to plug it in

hartnn (hartnn):

x1=2 y1=3 x2=-3 y2=0

OpenStudy (anonymous):

slope (m) formula: m=x2−x1/y2−y1 So plug in your numbers (2,3), (-3,0) -3-2/0-3 -5/-3 5/3 So five thirds

OpenStudy (anonymous):

Yup! I got that too! Thanks guys, :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

np but do u understand

OpenStudy (anonymous):

yes i do:)

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

your cheating

OpenStudy (anonymous):

No. He explained how.

OpenStudy (anonymous):

That is not cheating.

OpenStudy (anonymous):

na that don't look like that

OpenStudy (anonymous):

ok